Home‎ > ‎Reports‎ > ‎Crew103‎ > ‎

day12


Crew 103 Log Book for April 20, 2011

Photos of the day





Commander's Check-In Report

Anne-Sophie Kalbfleisch Reporting

Crew Physical Status:

Everybody is fine

Time Departed/Returned from EVA:

No EVA today. The weather was too bad

Brief Narrative of Field Mission Results:

Today we spent a lot of time filling the psychological reports and spent 1h30 for the psychological test from Ayako Ono, from University of Tohoku, for her Moon Mars Habitability Project. It was very windy so we weren't able to launch our hot-air balloon. We also continued our algae measurements on solid substrate and began a new experience with the liquid solution, including red LEDs in order to see their influence on their growth.

We were also very glad to detect some magnetic disturbances thanks to our self made radiotelescope. More information about the results will be given in a complete science report at the end of our stay.

EVA Data/Interpretations:

No EVA

Engineering/Hab Maintenance:

Nothing to mention

Plans for Tomorrow:

Today we couldn't go outside because of the bad weather and an EVA with the ATV's was too risky. We hope the weather will be better so that we could get back to the places where we took the soil samples (the battery of our GPS went off so we weren't able to note down the coordinates). We will also try to launch our hot-air balloon and continue our algae's measurements. We are very happy because they are growing well! We are just afraid of some traces of impurities in our samples. We hope they won't affect their growth.

Support Requested:

Nothing to mention



Commander's Report

Anne-Sophie Kalbfleisch Reporting

Today we spent a lot of time filling the psychological reports and spent 1h30 for the psychological test from Ayako Ono, from University of Tohoku, for her Moon Mars Habitability Project. During this workshop we had to note down which aspects could be improved, which aspects are important for a long term stay and so on. S. Viroux, as coordinator of this project, had to measure our blood pressure and our heartbeats.

It was very windy outside so we weren't able to launch our hot-air balloon. During the afternoon it even rained.

We also continued our algae measurements on solid substrate and began a new experience with the liquid solution, including red LEDs in order to see their influence on the algae growth. They are growing well and we are very happy about that, but we noticed some impurities on the solid substrate. We hope they won't affect their growth.

We have also detected some magnetic disturbances thanks to our self made radio-telescope. J. Offergeld and G. Van Roy will detail this results in their final science report. Tomorrow we will spend some time for the final reports and for the interpretation of our experimental results.


Engineering Report

Thibault Windels Reporting

Generator/Electricity:

Honey (Diesel Generator) (Used? Yes/No): yes

Wendy (Gas Generator) (Used? Yes/No): no

Battery Charge Level: 10,3 volts

Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):


Fuel Status:

Diesel (Full/0.75/0.5/0.25/Empty): 0,48

Propane Tank (Full/0.75/0.5/0.25/Empty): 0,36

Gasoline Tank (Full/0.75/0.5/0.25/Empty): 5 gallons approximately

Notes/Comments: /


Oil:

Kitty Oil Quantity (# of Quarts in storage): the previous commander told us not to check the oil quantity of Kitty because we have to turn the all generator off!

Honey Oil Quantity (# of Quarts in storage): /

ATV Oil Quantity (# of Quarts in storage): full

Car Oil Quantity (# of Quarts in storage): /


Water Status:

Outside Potable Water Tank Level (inches from bottom): 14,5

Trailer Potable Water Tank Level (inches from bottom): 7,5

Hab Potable Water Tank Level (inches from bottom): 11

Potable Water Meter Reading: 45280,8

In to GreenHab Meter Reading: /

Notes/Comments: /


GreenHab:

Condition of plants in Tank 1: /

Condition of plants in Tank 2: /

Condition of plants in Tank 3: /

Greenhab Notes/Comments: One of the two ventilators is on. All the thermostats are down to 50. One of the propane heaters is still putting out heat. the temperature inside is 87°F


Observatory:

Telescope Used? (Yes/No): no

Observatory Notes/Comments: there's no telescope


Transportation:

Hab Car (NOT RENTAL)

Used (yes/no): no

Oil Checked (yes/no): no

Tire Status: ok

Odometer at end of day: 124250,5

Notes/Comments on Hab Car: /


ATV 1 (Opportunity):

Used: no

Oil Checked: yes (we refilled the tank)

Fuel consumed : 0

Tire Status: ok

Comments: Sometimes it's difficult to change the gear.


ATV 2 (Spirit):

Used: no

Oil Checked: yes

Fuel Consumed: 0

Tire Status: ok

Comments: We have to start manually the ATV because the electric button doesn't work anymore.


ATV 3 (Viking I):

Used: no

Oil Checked: yes

Fuel Consumed: 0

Tire Status: ok

Comments: /


ATV 4 (Viking II):

Used: no

Oil Checked: no

Fuel Consumed: 0

Tire Status:

Comments: does not work


Heating and Ventilation:

Thermostat setting upstairs (Farenheit): 65

Actual temperature upstairs (Farenheit): 77

Thermostat setting downstairs (Farenheit): /

Actual temperature downstairs (Farenheit): 75


Computers/Networking Infrastructure:

Notes/comments on internet/computers?: internet is still running well

General Engineering Notes/Comments: /


Science Report (Biology)

Wouter de Groot Reporting

Today we ended the first part of our experience in aqueous environment. This first part was the study of their development under blue LED so we made our last measurements of them, we dissolved the solution to get the same concentration as the beginning of the first part and we replace one blue LED by a red one. This replacement is made to study the influence of a mixed spectrum compared to a pure one. If the evolution of the optical density is accelerated by this way it will show that the mixed character of the spectrum improves algae development. Then we restarted the measurements, following the same procedure as usual. We will send you a picture of the new assembly.

I don't know if you want to know the exact results of the experiences, I was going to send it in my general LED report. If you want to see them before, please ask me!



Chef's Report

Gregory Van Roy Reporting

1a) Today was

A cooking day

1b) Was today a special day celebrated at the main meal with special food or activity?

Nothing special today

2)  List the foods served at today's main meal (usually the evening meal), giving full names of each.

Beerbreak Navy Bean Soup, Tortilla with vegetables, tomato sauce and salmon and pomegranate drink mix

3) List any main meal foods not finished by the crew (leftovers stored for later use or discarded)

Tortilla : 2 packs left

Tomato sauce : 1/4 of the pack left

4)  Recipes for anything more complicated than following package

directions exactly, or rehydrating.

None

5)  Lessons learned  (if any)

None

6) Comments/questions for Kim and Jean, the food study investigators

We saw a lot of tofu in the rack, I think nobody use it because it's not really tasty and it's complicated to cook it... We will try this evening to do something good with chocolate but recipes seem being difficult!